# The Mean-Field Dynamics of Transformers

### v2

@@ Abstract.
Transformers, introduced by Vaswani et al. [vaswani2017attention], have become the dominant architecture in modern machine learning, powering large language models and state-of-the-art systems across modalities. Their key novelty lies in the attention mechanism, a data-dependent interaction between components of an input sequence. This mechanism allows each token—a vector representing a word, image patch, or general embedding—to update its representation by attending to all other tokens in the sequence.
From a mathematical standpoint, the attention mechanism can be viewed as defining a pairwise interaction between tokens. Because neural networks act through an iterative composition of layers, their evolution may be interpreted as a discrete-time dynamical system, and in the continuous-time limit, as a nonlinear flow. This viewpoint underlies the theory of neural ordinary differential equations (neural ODEs) [weinan2017proposal, haber2017stable, chen2018neural, li2018maximum, weinan2019mean, ruiz2023neural], and provides a natural mathematical lens through which to analyze deep architectures. From this perspective, a Transformer may be seen as an interacting particle system, in which each particle follows a velocity field depending on the empirical distribution of all others. This situates Transformers within the broad mathematical framework of mean-field dynamics.
@@
The Transformer architecture, introduced by Vaswani et al. [vaswani2017attention], is a modular neural network designed to process sequences of variable length. In contrast with recurrent and convolutional networks, whose structure enforces local or sequential dependencies, Transformers rely on a global interaction mechanism known as self-attention or simply attention. This mechanism allows every element of an input sequence—called a token—to interact with every other token in a data-dependent way.
@@
$X_{i}$ $\longmapsto X+\mathrm{MLP}(X_{i})\,$,
for $i=1,\ldots,n$. In particular, all the tokens interact in the Attention layer but not in the MLP layer. Normalization steps are inserted before or after each layer depending on the implementation; see Section 6.2 for some examples. These residual connections, together with layer normalization [ba2016layer, zhang2019root], play a crucial stabilizing role, analogous to skip connections in residual networks [he2016identity, haber2017stable]. We will later see that such normalization can be idealized as a projection onto the unit sphere, a simplification that preserves essential dynamical features of trained models.
@@
Multi-headed attention and MLP blocks.
In practice, the attention computation is distributed across several heads, indexed by $h=1,\dots,H$, each with its own triplet $(Q_{h},K_{h},V_{h})$. The resulting outputs are concatenated and linearly recombined. Multi-headed attention increases expressiveness and parallelism, and may be viewed as sampling several interaction kernels in parallel. Following the attention block, a feed-forward network (usually a two-layer multilayer perceptron) acts independently on each token:
@@
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function [hendrycks2016gaussian] applied to each entry of the vector $W_{1}x+b_{1}$; a simple example to keep in mind is $\sigma(t)=\tanh(t)$ but many variations exist [kunc2024decadesactivationscomprehensivesurvey]. Together, the attention and MLP components define the local and global dynamics of the model.
@@
Transformers were first developed for sequence-to-sequence tasks such as translation and therefore include an encoder–decoder structure. The encoder transforms an input sequence into a latent representation through successive attention and MLP layers; the decoder then generates outputs by combining this representation with a causal (i.e., temporally restricted) version of attention. Variants such as BERT [devlin2019bert] use only the encoder, whereas GPT-type models employ the decoder alone with causal masking to enforce autoregressive behavior. Throughout this paper we focus on the encoder-type architecture and analyze the attention dynamics without causal masking, though we note that causal attention exhibits closely related phenomena [karagodin2024clustering].
@@
Normalization layers play a crucial role in stabilizing training and controlling the geometry of representations. The two main variants are post-layer normalization (which projects tokens back onto the sphere after applying attention) and pre-layer normalization (which projects tokens onto the sphere before attention). Pre-layer normalization is used in leading models such as GPT [radford2019gpt] and LLaMA [touvron2023llama]. While this paper focuses on post-layer normalization, the mathematical results are not crucially dependent on that choice. In fact, we describe in Section 6.2 how several popular normalization rules can be understood in a common framework and all lead to some clustering, albeit at different speeds.
@@
Modern variants of the Transformer introduce additional components: hierarchical structures for vision tasks [liu2021swin], sparse or long-range attention [beltagy2020longformer, child2019generating], low-rank adaptation mechanisms [hu2022lora], and mixture-of-experts [shazeer2017outrageously, boixadsera2025powerfinegrainedexpertsgranularity] architectures that route tokens through specialized subnetworks. These extensions have led to dramatic empirical gains but lie outside the scope of this paper. Our goal is to isolate and analyze the mathematical core of the architecture, the attention mechanism itself, and to understand how it regulates token dynamics.
@@
$X_{k+1}=X_{k}+F_{k}(X_{k}),$
where $X_{k}\in\mathbb{R}^{n\times d}$ denotes the matrix of token embeddings at layer $k$, and $F$ encodes the combination of attention, feed-forward, and normalization operations. This recursive structure naturally suggests a discrete-time dynamical system. Following the analogy with residual neural networks [he2016identity, haber2017stable, chen2018neural, weinan2017proposal], we interpret the layer index as a discretized time variable, and pass to the continuous-time limit
$\dot{X}_{t}=F_{t}(X_{t}).$
@@
$\dot{x}_{i}(t)=\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle x_{i}(t),x_{j}(t)\rangle}\,x_{j}(t),$ (USA)
@@
$\partial_{t}\mu_{t}+\nabla\!\cdot(\mu_{t}\,v_{t}[\mu_{t}])=0,\qquad v_{t}(x)=\int e^{\beta\langle x,y\rangle}y\,\mathrm{d}\mu_{t}(y).$ (2)
@@
This continuum formulation reveals an important structural distinction between the normalized and unnormalized models. For (USA), the partial differential equation (PDE) (2) is the Wasserstein gradient flow [ambrosio2005gradient, CheNilRig25] of the interaction energy
@@
For (SA), it also corresponds to a gradient flow over the space of probability measures, albeit with respect to a different (Hessian) metric [geshkovski2025mathematical, Li21]. In the large-$\beta$ regime, these two dynamics recover familiar PDEs at leading order. Indeed the continuity equation associated with (SA) converges formally to a reverse heat equation. This anti-diffusive limit foreshadows the clustering behavior described in the next section. In contrast, for (USA) with appropriate time rescaling, it converges to a porous medium equation; see [bruno2025multiscale].
@@
the classical (homogeneous) Kuramoto model [kuramoto1975self, acebron2005kuramoto], originally introduced to study synchronization of coupled oscillators. It is known that for such dynamics, trajectories synchronize: for almost every initialization $(\theta_{1}(0),\ldots,\theta_{n}(0))\in\mathbb{T}^{n}$, one has $|\theta_{i}(t)-\theta_{j}(t)|\to 0$ as $t\to\infty$ [taylor2012there].
@@
Figure 1. Histograms of pairwise inner products $\{\langle x_{i}(t),x_{j}(t)\rangle\}_{i\neq j}$ at layers $t=0$, 55, 3030, and $48$ in the pre-trained ALBERT XLarge v2 model [lanalbert] available on Huggingface. The progressive concentration of mass near 11 across layers illustrates the emergence of clustering in token embeddings. Figure reproduced from [geshkovski2025mathematical].
@@
We now recall a general result guaranteeing convergence of the dynamics (SA) to a clustered state for all temperatures $\beta\geq 0$ and all ambient dimensions $d\geq 3$. The result originates in the study of synchronization on spheres, a line of work initiated by [markdahl2017almost] and recently refined in [criscitiello2024synchronization]. It asserts that for a large class of smooth interaction laws depending only on pairwise inner products, the only asymptotically stable equilibria correspond to complete synchronization.
Theorem 1 ([markdahl2017almost, criscitiello2024synchronization, geshkovski2025mathematical]).
The following holds for both (SA) and (USA) dynamics with $n\geq 2$ particles in dimension $d\geq 3$ and any $\beta\geq 0$. For almost every initial condition $(x_{1}(0),\dots,x_{n}(0))\in(\mathbb{S}^{d-1})^{n}$, the trajectories exist globally and converge to a fully synchronized configuration:
@@
The result follows from the general theory of consensus dynamics on compact manifolds [markdahl2017almost]. Since the dynamics (SA) are a smooth gradient flow of an analytic energy functional, the classical theorem of Łojasiewicz [lojasiewicz1963propriete] guarantees that every trajectory converges to a stationary point. A detailed analysis of the critical set shows that, except for the fully synchronized configuration, all other stationary points are saddles by exhibiting an escape direction. This result also holds for weighted tokens that evolve according to the Wasserstein-Fisher-Rao gradient flow whereby particles evolve and are also dynamically reweighted [chen2025clustering]. By the center-stable manifold theorem [shub2013global, Thm. III.7], the set of initial conditions whose trajectories converge to such saddles is contained in a countable union of lower-dimensional manifolds and thus has measure zero. Consequently, almost every initialization leads to convergence toward the synchronized equilibrium. This argument, first established in [markdahl2017almost] and subsequently refined in [criscitiello2024synchronization], provides a complete characterization of the long-time behavior of the (USA) dynamics; see [geshkovski2025mathematical] for an extension to the (SA) dynamics.
@@
The case $d=2$ (the circle) is not covered by Theorem 1, although it corresponds to the Kuramoto model ($\beta=0$) of Section 3.3 for which synchronization was proved in [taylor2012there]. Indeed, the proof of Theorem 1 fails when $d=2$. This gap was recently closed in [andrew25], which proves clustering for all $\beta>-0.16$ in the case $d=2$ for both (USA) and (SA).
@@
For the Kuramoto model (the case $d=2$, $\beta=0$), Morales and Poyato [morales2022trend] established exponential rates of convergence of $\mu_{t}$ to a Dirac point mass. In higher dimensions and for general $\beta>0$, the following theorem extends this result to self-attention dynamics.
Theorem 5 ([chen2025quantitative]).
@@
This result provides a quantitative convergence rate for attention dynamics in the mean-field regime, complementing the qualitative clustering theorems for finite particles. The argument extends that of [morales2022trend] to arbitrary dimensions but it is limited to small $\beta$. Indeed, an upper bound on $\beta$ is necessary: Example 2.6 in [chen2025quantitative] constructs an initialization on the circle showing that for large $\beta$ the mean-field dynamics may converge to multiple clusters. Such behavior suggests that large values of $\beta$ create an increasingly complex energy landscape in which several metastable states may appear as discussed in the next section.
@@
The previous section established that attention dynamics lead to clustering, either for finitely many particles or at the mean-field level. While these results characterize the asymptotic limit in which all particles collapse to a single cluster, empirical evidence (Figure 1) and numerical simulations (white regions in Figure 3) show that in practice, multiple clusters are typically observed. Such multi-cluster configurations are in fact desirable: when Transformers are viewed as measure-to-measure maps, they substantially enhance expressivity compared to the degenerate single-cluster limit. The early formation of these clusters has been rigorously analyzed in [bruno2025emergence], which, in the mean-field limit, describes how small perturbations around the uniform initialization amplify into a structured, periodic arrangement whose number of clusters depends on the temperature parameter $\beta$. A complementary analysis in [bruno2025multiscale] extends this picture to general parameter matrices and identifies multiple dynamical phases, including a slow phase associated with metastability, though it reduces to the models considered here when these matrices are the identity. Here, we focus our exposition on the long-time evolution of these clusters—specifically, on their slow coalescence and eventual collapse, a regime that reveals the metastable character of attention dynamics.
@@
The analysis in [geshkovski2024dynamic] establishes the existence of exponentially long-lived metastable states. Starting from a well-separated initial configuration consisting of $k$ sets of tokens that are close together relative to other sets, the tokens first collapse within each set, forming $k$ tight groups during the time interval $0,T_{1}]$, where $T_{1}\sim\beta$. These clusters then persist on an interval $T_{1},T_{2}]$ where $\log T_{2}\sim\beta$ before successive merging events occur. Geometrically, the flow of tokens remains near a manifold of $k$-cluster configurations, where the energy gradient is exponentially small and the motion of clusters is correspondingly slow.
This metastable behavior can be understood from the energy perspective. A configuration with $k>1$ well-separated clusters corresponds to a nearly stationary point of $\mathcal{E}_{\beta}$ where $\|\nabla\mathcal{E}_{\beta}\|$ is exponentially small. Consequently, trajectories evolve extremely slowly in such regions of the energy landscape. This fits into the slow-motion framework of Otto and Reznikoff [otto2007slow], which states that for a gradient flow $\dot{X}=-\nabla\mathcal{E}(X)$, if $\|\nabla\mathcal{E}\|\leq\delta\ll 1$ on a manifold $\mathcal{M}$ and $\mathcal{E}$ satisfies a Polyak–Łojasiewicz-type inequality near $\mathcal{M}$, then trajectories remain trapped near $\mathcal{M}$ for times of order $\delta^{-1}$. Here $\delta\sim e^{-c\beta}$, explaining the exponentially long metastable time scale.
@@
The energy functional $\mathcal{E}_{\beta}$ introduced in (3) admits a hierarchy of saddle points of increasing energy, connected by heteroclinic orbits that describe the gradual merging of clusters. After a suitable time rescaling, the energy evolves through long plateaux corresponding to metastable phases, separated by abrupt jumps each time two clusters coalesce. These saddle-to-saddle transitions become sharpest in the limit of the gradient flow of $\mathcal{E}_{\beta}$ when $\beta\to\infty$ and time is properly rescaled. In this regime, the configuration remains effectively frozen until it moves abruptly to the next saddle in the hierarchy by merging two clusters.
Figure 2. Energy $\mathcal{E}_{\beta}(t)$ along a metastable trajectory in the low-temperature limit $\beta\to\infty$. The energy remains constant over long plateaus corresponding to metastable multi-cluster configurations and increases sharply when clusters merge, forming a staircase profile. Each jump corresponds to a transition between successive saddle points of the energy landscape. Figure reproduced from [geshkovski2024dynamic].
@@
Following [geshkovski2024dynamic], Bruno, Pasqualotto, and Agazzi [bruno2025multiscale] provide a complete multiscale dynamical picture describing the evolution from such an initialization to the final single-cluster state in the limit $\beta\to\infty$. In particular, their analysis identifies a final pairing phase governed by hardmax-like dynamics [hardmax], during which the two closest clusters merge. This corresponds to a transition from one saddle of $\mathcal{E}_{\beta}$ to the next higher-energy saddle.
To build intuition, notice that as $\beta\to\infty$ the softmax velocity field in (SA) converges [^1] Formally, the limiting rule is $x_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\!\left(\mathop{\mathrm{arg\,max}}_{x_{j}(t)}\langle x_{i}(t),x_{j}(t)\rangle\right)={\mathbf{P}}^{\perp}_{x_{i}(t)}x_{i}(t)=0\,.$ Following [hardmax], and to avoid this triviality, we forbid tokens from attending to themselves. to a hard $\mathop{\mathrm{arg\,max}}$ rule:
@@
The dynamics (SA) are closely related to a continuous-time analogue of the classical Mean-Shift clustering algorithm. In its original form, the Mean-Shift algorithm defines clusters as basins of attraction of the modes of a kernel density estimator (KDE) [fukunaga1975estimation]. More precisely, given points $x_{1}(0),\ldots,x_{n}(0)\in\mathbb{R}^{d}$ drawn independently from a density $p$, and a kernel $K(\cdot)$ on $\mathbb{R}^{d}$ (typically Gaussian), recall that the KDE of $p$ is given by $\hat{p}=K*\mu_{0}$, the convolution of $K$ with the empirical measure $\mu_{0}$ of the $x_{i}$; see, e.g., [tsybakov2008nonparametric, Chapter 1]. The following gradient-flow dynamics move each point toward the nearest mode of $K*\mu_{0}$:
@@
This algorithm, along with suitable time discretizations, can be analyzed using tools from optimization and statistics. Notably, [arias2016estimation] establish consistency of the estimated gradient lines and show that fixed-KDE Mean-Shift recovers the modal structure of the underlying density under classical smoothness assumptions.
A modification proposed by Cheng [cheng1995mean], often called blurring Mean-Shift, recomputes the KDE at every iteration using the updated points. This leads to the mean-field dynamics
@@
This analogy suggests tokens cluster into $M$ groups, where $M$ is the number of modes of the KDE $\hat{p}$. Using Edgeworth expansions and the Kac–Rice formula, we can show that $\mathbb{E}[M]$ is of order $\sqrt{\beta\log\beta}$ as $n\to\infty$, in the regime $n^{c}\leq\beta\leq n^{2-c}$ (with arbitrarily small fixed $c$), when $p$ is a Gaussian density on the real line [geshkovski2025numbermodesgaussiankernel]. Numerical simulations indicate that this order of magnitude should also hold on the circle, at least up to logarithmic factors. Recall, however, that this multiple-cluster state is metastable and will eventually collapse to a single one according to the results of Section 4.
While this approach is currently limited to the one-dimensional case, it gives an approach to compute the number of metastable states in self-attention dynamics. An alternative approach to understand the first metastable state was proposed in [bruno2025emergence] by linearizing (2) around the uniform distribution. This different approach also points to a metastable state that contains $\Theta(\sqrt{\beta})$ clusters.
@@
This equiangular model was initially introduced in [geshkovski2025mathematical] to study the evolution of (near-)orthogonal token initialization in attention dynamics and was later used in [cowsik2024geometric] and [giorlandino2025failuremodesdeeptransformers] to examine the role of large random weight matrices using formal heuristic calculations.
@@
with initial condition $\rho(0)=\rho_{0}$. Linearizing near the synchronized state $\rho=1$ and setting $\varepsilon(t)=1-\rho(t)$, we obtain
@@
Figure 3. Phase transition diagrams for randomly initialized particles following (SA). Each panel shows the probability that $\langle x_{i}(t),x_{j}(t)\rangle\geq 0.999$ on a fine grid of values for the pair $(t,\beta)$ for $n=32$ tokens. As the dimension $d$ increases from $d=8$ (left), to $d=128$ (middle) to $d=1,024$ (right), the transition curve sharpens and aligns with the theoretical prediction from the equiangular model. Reproduced from [geshkovski2025mathematical].

### v3

@@ Abstract.
Transformers, introduced by Vaswani et al. [67], have become the dominant architecture in modern machine learning, powering large language models and state-of-the-art systems across modalities. Their key novelty lies in the attention mechanism, a data-dependent interaction between components of an input sequence. This mechanism allows each token—a vector representing a word, image patch, or general embedding—to update its representation by attending to all other tokens in the sequence.
From a mathematical standpoint, the attention mechanism can be viewed as defining a pairwise interaction between tokens. Because neural networks act through an iterative composition of layers, their evolution may be interpreted as a discrete-time dynamical system, and in the continuous-time limit, as a nonlinear flow. This viewpoint underlies the theory of neural ordinary differential equations (neural ODEs) [26, 35, 15, 46, 25, 59], and provides a natural mathematical lens through which to analyze deep architectures. From this perspective, a Transformer may be seen as an interacting particle system, in which each particle follows a velocity field depending on the empirical distribution of all others. This situates Transformers within the broad mathematical framework of mean-field dynamics.
@@
The Transformer architecture, introduced by Vaswani et al. [67], is a modular neural network designed to process sequences of variable length. In contrast with recurrent and convolutional networks, whose structure enforces local or sequential dependencies, Transformers rely on a global interaction mechanism known as self-attention or simply attention. This mechanism allows every element of an input sequence—called a token—to interact with every other token in a data-dependent way.
@@
$X_{i}$ $\longmapsto X_{i}+\mathrm{MLP}(X_{i})\,$,
for $i=1,\ldots,n$. In particular, all the tokens interact in the Attention layer but not in the MLP layer. Normalization steps are inserted before or after each layer depending on the implementation; see Section 6.2 for some examples. These residual connections, together with layer normalization [5, 68], play a crucial stabilizing role, analogous to skip connections in residual networks [36, 35]. We will later see that such normalization can be idealized as a projection onto the unit sphere, a simplification that preserves essential dynamical features of trained models.
@@
In practice, the attention computation is distributed across several heads, indexed by $h=1,\dots,H$, each with its own triplet $(Q_{h},K_{h},V_{h})$. The resulting outputs are concatenated and linearly recombined. Multi-headed attention increases expressiveness and parallelism, and may be viewed as sampling several interaction kernels in parallel.
MLP blocks.
Following the attention block, a feed-forward network (usually a two-layer multilayer perceptron) acts independently on each token:
@@
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function [37] applied to each entry of the vector $W_{1}x+b_{1}$; a simple example to keep in mind is $\sigma(t)=\tanh(t)$ but many variations exist [42]. Together, the attention and MLP components define the local and global dynamics of the model.
@@
Transformers were first developed for sequence-to-sequence tasks such as translation and therefore include an encoder–decoder structure. The encoder transforms an input sequence into a latent representation through successive attention and MLP layers; the decoder then generates outputs by combining this representation with a causal (i.e., temporally restricted) version of attention. Variants such as BERT [24] use only the encoder, whereas GPT-type models employ the decoder alone with causal masking to enforce autoregressive behavior. Throughout this paper we focus on the encoder-type architecture and analyze the attention dynamics without causal masking, though we note that causal attention exhibits closely related phenomena [40].
@@
Normalization layers play a crucial role in stabilizing training and controlling the geometry of representations. The two main variants are post-layer normalization (which projects tokens back onto the sphere after applying attention) and pre-layer normalization (which projects tokens onto the sphere before attention). Pre-layer normalization is used in leading models such as GPT [58] and LLaMA [65]. While this paper focuses on post-layer normalization, the mathematical results are not crucially dependent on that choice. In fact, we describe in Section 6.2 how several popular normalization rules can be understood in a common framework and all lead to some clustering, albeit at different speeds.
@@
Modern variants of the Transformer introduce additional components: hierarchical structures for vision tasks [48], sparse or long-range attention [8, 20], low-rank adaptation mechanisms [38], and mixture-of-experts [62, 10] architectures that route tokens through specialized subnetworks. These extensions have led to dramatic empirical gains but lie outside the scope of this paper. Our goal is to isolate and analyze the mathematical core of the architecture, the attention mechanism itself, and to understand how it regulates token dynamics.
@@
$\mathbf{X}_{k+1}=\mathbf{X}_{k}+F_{k}(\mathbf{X}_{k}),$
where $\mathbf{X}_{k}\in\mathbb{R}^{n\times d}$ denotes the matrix of token embeddings at layer $k$, and $F$ encodes the combination of attention, feed-forward, and normalization operations. This recursive structure naturally suggests a discrete-time dynamical system. Following the analogy with residual neural networks [36, 35, 15, 26], we interpret the layer index as a discretized time variable, and pass to the continuous-time limit
$\dot{\mathbf{X}}_{t}=F_{t}(\mathbf{X}_{t}).$
@@
$\dot{x}_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\!\left(\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle x_{i}(t),x_{j}(t)\rangle}\,x_{j}(t)\right),$ (USA)
@@
$\partial_{t}\mu_{t}+\nabla\!\cdot(\mu_{t}\,v_{t}[\mu_{t}])=0,\qquad v_{t}(x)={\mathbf{P}}^{\perp}_{x}\!\int e^{\beta\langle x,y\rangle}y\,\mathrm{d}\mu_{t}(y).$ (2)
@@
This continuum formulation reveals an important structural distinction between the normalized and unnormalized models. For (USA), the partial differential equation (PDE) (2) is the Wasserstein gradient flow [3, 19] of the interaction energy
@@
For (SA), it also corresponds to a gradient flow over the space of probability measures, albeit with respect to a different (Hessian) metric [32, 47]. In the large-$\beta$ regime, these two dynamics recover familiar PDEs at leading order. Indeed the continuity equation associated with (SA) converges formally to a reverse heat equation. This anti-diffusive limit foreshadows the clustering behavior described in the next section. In contrast, for (USA) with appropriate time rescaling, it converges to a porous medium equation; see [11].
@@
the classical (homogeneous) Kuramoto model [43, 1], originally introduced to study synchronization of coupled oscillators. It is known that for such dynamics, trajectories synchronize: for almost every initialization $(\theta_{1}(0),\ldots,\theta_{n}(0))\in\mathbb{T}^{n}$, one has $|\theta_{i}(t)-\theta_{j}(t)|\to 0$ as $t\to\infty$ [64].
@@
Figure 1. Histograms of pairwise inner products $\{\langle x_{i}(t),x_{j}(t)\rangle\}_{i\neq j}$ at layers $t=0$, 55, 3030, and $48$ in the pre-trained ALBERT XLarge v2 model [44] available on Huggingface. The progressive concentration of mass near 11 across layers illustrates the emergence of clustering in token embeddings. Figure reproduced from [32].
@@
We now recall a general result guaranteeing convergence of the dynamics (SA) to a clustered state for all temperatures $\beta\geq 0$ and all ambient dimensions $d\geq 3$. The result originates in the study of synchronization on spheres, a line of work initiated by [51] and recently refined in [22]. It asserts that for a large class of smooth interaction laws depending only on pairwise inner products, the only asymptotically stable equilibria correspond to all particles being clustered in a single location (complete synchronization).
Theorem 1 ([51, 22, 32]).
The following holds for both (SA) and (USA) dynamics with $n\geq 2$ particles in dimension $d\geq 3$ and any $\beta\geq 0$. For almost every initial condition $(x_{1}(0),\dots,x_{n}(0))\in(\mathbb{S}^{d-1})^{n}$, the trajectories exist globally and converge to a clustered configuration:
@@
The result follows from the general theory of consensus dynamics on compact manifolds [51]. Since the dynamics (SA) are a smooth gradient flow of an analytic energy functional, the classical theorem of Łojasiewicz [49] guarantees that every trajectory converges to a stationary point. A detailed analysis of the critical set shows that, except for the clustered configuration, all other stationary points are saddles by exhibiting an escape direction. This result also holds for weighted tokens that evolve according to the Wasserstein-Fisher-Rao gradient flow whereby particles evolve and are also dynamically reweighted [17]. By the center-stable manifold theorem [63, Thm. III.7], the set of initial conditions whose trajectories converge to such saddles is contained in a countable union of lower-dimensional manifolds and thus has measure zero. Consequently, almost every initialization leads to convergence toward the clustered equilibrium. This argument, first established in [51] and subsequently refined in [22], provides a complete characterization of the long-time behavior of the (USA) dynamics; see [32] for an extension to the (SA) dynamics.
@@
The case $d=2$ (the circle) is not covered by Theorem 1, although it corresponds to the Kuramoto model ($\beta=0$) of Section 3.3 for which synchronization/clustering was proved in [64]. Indeed, the proof of Theorem 1 fails when $d=2$. This gap was recently closed in [56], which proves clustering for all $\beta>-0.16$ in the case $d=2$ for both (USA) and (SA).
@@
For the Kuramoto model (the case $d=2$, $\beta=0$), Morales and Poyato [52] established exponential rates of convergence of $\mu_{t}$ to a Dirac point mass. In higher dimensions and for general $\beta>0$, the following theorem extends this result to self-attention dynamics.
Theorem 5 ([16]).
@@
This result provides a quantitative convergence rate for attention dynamics in the mean-field regime, complementing the qualitative clustering theorems for finite particles. The argument extends that of [52] to arbitrary dimensions but it is limited to small $\beta$. Indeed, an upper bound on $\beta$ is necessary: Example 2.6 in [16] constructs an initialization on the circle showing that for large $\beta$ the mean-field dynamics may converge to multiple clusters. Such behavior suggests that large values of $\beta$ create an increasingly complex energy landscape in which several metastable states may appear as discussed in the next section.
@@
The previous section established that attention dynamics lead to clustering, either for finitely many particles or at the mean-field level. While these results characterize the asymptotic limit in which all particles collapse to a single cluster, empirical evidence (Figure 1) and numerical simulations (white regions in Figure 3) show that in practice, multiple clusters are typically observed. Such multi-cluster configurations are in fact desirable: when Transformers are viewed as measure-to-measure maps, they substantially enhance expressivity compared to the degenerate single-cluster limit. The early formation of these clusters has been rigorously analyzed in [12], which, in the mean-field limit, describes how small perturbations around the uniform initialization amplify into a structured, periodic arrangement whose number of clusters depends on the temperature parameter $\beta$. A complementary analysis in [11] extends this picture to general parameter matrices and identifies multiple dynamical phases, including a slow phase associated with metastability, though it reduces to the models considered here when these matrices are the identity. Here, we focus our exposition on the long-time evolution of these clusters—specifically, on their slow coalescence and eventual collapse, a regime that reveals the metastable character of attention dynamics.
@@
The analysis in [30] establishes the existence of exponentially long-lived metastable states. Starting from a well-separated initial configuration consisting of $k$ sets of tokens that are close together relative to other sets, the tokens first collapse within each set, forming $k$ tight groups during the time interval $0,T_{1}]$, where $T_{1}\sim\beta$. These clusters then persist on an interval $T_{1},T_{2}]$ where $\log T_{2}\sim\beta$ before successive merging events occur. Geometrically, the flow of tokens remains near a manifold of $k$-cluster configurations, where the energy gradient is exponentially small and the motion of clusters is correspondingly slow.
This metastable behavior can be understood from the energy perspective. A configuration with $k>1$ well-separated clusters corresponds to a nearly stationary point of $\mathcal{E}_{\beta}$ where $\|\nabla\mathcal{E}_{\beta}\|$ is exponentially small. Consequently, trajectories evolve extremely slowly in such regions of the energy landscape. This fits into the slow-motion framework of Otto and Reznikoff [55], which states that for a gradient flow $\dot{X}=-\nabla\mathcal{E}(X)$, if $\|\nabla\mathcal{E}\|\leq\delta\ll 1$ on a manifold $\mathcal{M}$ and $\mathcal{E}$ satisfies a Polyak–Łojasiewicz-type inequality near $\mathcal{M}$, then trajectories remain trapped near $\mathcal{M}$ for times of order $\delta^{-1}$. Here $\delta\sim e^{-c\beta}$, explaining the exponentially long metastable time scale.
@@
The energy functional $\mathcal{E}_{\beta}$ introduced in (3) admits a hierarchy of saddle points of increasing energy, connected by heteroclinic orbits that describe the gradual merging of clusters. After a suitable time rescaling, the energy evolves through long plateaus corresponding to metastable phases, separated by abrupt jumps each time two clusters coalesce. These saddle-to-saddle transitions become sharpest in the limit of the gradient flow of $\mathcal{E}_{\beta}$ when $\beta\to\infty$ and time is properly rescaled. In this regime, the configuration remains effectively frozen until it moves abruptly to the next saddle in the hierarchy by merging two clusters.
Figure 2. Energy $\mathcal{E}_{\beta}(t)$ along a metastable trajectory in the low-temperature limit $\beta\to\infty$. The energy remains constant over long plateaus corresponding to metastable multi-cluster configurations and increases sharply when clusters merge, forming a staircase profile. Each jump corresponds to a transition between successive saddle points of the energy landscape. Figure reproduced from [30].
@@
Following [30], Bruno, Pasqualotto, and Agazzi [11] provide a complete multiscale dynamical picture describing the evolution from such an initialization to the final single-cluster state in the limit $\beta\to\infty$. In particular, their analysis identifies a final pairing phase governed by hardmax-like dynamics [2], during which the two closest clusters merge. This corresponds to a transition from one saddle of $\mathcal{E}_{\beta}$ to the next higher-energy saddle.
To build intuition, notice that as $\beta\to\infty$ the softmax velocity field in (SA) converges [^1] Formally, the limiting rule is $x_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\!\left(\mathop{\mathrm{arg\,max}}_{x_{j}(t)}\langle x_{i}(t),x_{j}(t)\rangle\right)={\mathbf{P}}^{\perp}_{x_{i}(t)}x_{i}(t)=0\,.$ Following [2], and to avoid this triviality, we forbid tokens from attending to themselves. to a hard $\mathop{\mathrm{arg\,max}}$ rule:
@@
The dynamics (SA) are closely related to a continuous-time analogue of the classical Mean-Shift clustering algorithm. In its original form, the Mean-Shift algorithm defines clusters as basins of attraction of the modes of a kernel density estimator (KDE) [27]. More precisely, given points $x_{1}(0),\ldots,x_{n}(0)\in\mathbb{R}^{d}$ drawn independently from a density $p$, and a kernel $K(\cdot)$ on $\mathbb{R}^{d}$ (typically Gaussian), recall that the KDE of $p$ is given by $\hat{p}=K*\mu_{0}$, the convolution of $K$ with the empirical measure $\mu_{0}$ of the $x_{i}$; see, e.g., [66, Chapter 1]. The following gradient-flow dynamics move each point toward the nearest mode of $K*\mu_{0}$:
@@
This algorithm, along with suitable time discretizations, can be analyzed using tools from optimization and statistics. Notably, [4] establish consistency of the estimated gradient lines and show that fixed-KDE Mean-Shift recovers the modal structure of the underlying density under classical smoothness assumptions.
A modification proposed by Cheng [18], often called blurring Mean-Shift, recomputes the KDE at every iteration using the updated points. This leads to the mean-field dynamics
@@
This analogy suggests tokens cluster into $M$ groups, where $M$ is the number of modes of the KDE $\hat{p}$. Using Edgeworth expansions and the Kac–Rice formula, we can show that $\mathbb{E}[M]$ is of order $\sqrt{\beta\log\beta}$ as $n\to\infty$, in the regime $n^{c}\leq\beta\leq n^{2-c}$ (with arbitrarily small fixed $c$), when $p$ is a Gaussian density on the real line [33]. Numerical simulations indicate that this order of magnitude should also hold on the circle, at least up to logarithmic factors. Recall, however, that this multiple-cluster state is metastable and will eventually collapse to a single one according to the results of Section 4.
While this approach is currently limited to the one-dimensional case, it gives an approach to compute the number of metastable states in self-attention dynamics. An alternative approach to understand the first metastable state was proposed in [12] by linearizing (2) around the uniform distribution. This different approach also points to a metastable state that contains $\Theta(\sqrt{\beta})$ clusters.
@@
This equiangular model was initially introduced in [32] to study the evolution of (near-)orthogonal token initialization in attention dynamics and was later used in [21] and [34] to examine the role of large random weight matrices using formal heuristic calculations.
@@
with initial condition $\rho(0)=\rho_{0}$. Linearizing near the clustered state $\rho=1$ and setting $\varepsilon(t)=1-\rho(t)$, we obtain
@@
Figure 3. Phase transition diagrams for randomly initialized particles following (SA). Each panel shows the probability that $\langle x_{i}(t),x_{j}(t)\rangle\geq 0.999$ on a fine grid of values for the pair $(t,\beta)$ for $n=32$ tokens. As the dimension $d$ increases from $d=8$ (left), to $d=128$ (middle) to $d=1,024$ (right), the transition curve sharpens and aligns with the theoretical prediction from the equiangular model. Reproduced from [32].